# Stable electron-irradiated [1-13C]alanine radicals for metabolic imaging with Dynamic Nuclear Polarisation Short Title: Electron-irradiated radicals for DNP

### v2

@@ Abstract
Dissolution Dynamic Nuclear Polarisation (dDNP) increases the sensitivity of magnetic resonance experiments by $10^{4}-$fold, permitting isotopically-labelled molecules to be transiently visible in MRI scans. dDNP mechanistically takes place at $\sim 1$ K, and requires unpaired electrons and microwaves. These electrons are usually chemical radicals, requiring removal by filtration prior to injection into humans. Alternative sources, such as UV irradiation, generate lower polarisation and require cryogenic transport. We present ultra-high-dose-rate electron irradiation as a novel alternative for generating non-persistent radicals in alanine/glycerol mixtures. These are stable for months at room temperature, quench spontaneously upon dissolution, are present in dose-dependent concentrations, and generate comparable nuclear polarisation (17%) to trityl radicals used clinically (19%) through a novel mechanism we believe to involve partial ordering, and electron-electron interactions. Owing to the large radiation doses required, this process is sterilising, permits imaging of alanine metabolism in vivo in the rat kidney, and may aid clinically translating dDNP.
@@ 1 Introduction
Hyperpolarised Magnetic Resonance Imaging (HP MRI) is a molecular imaging technique that is predominantly used to monitor metabolic fluxes in real-time by transiently making a labelled molecular probe visible to magnetic resonance experiments such as MRI. Over the past two decades, its pre-clinical use has spanned multiple fields of basic scientific and medical research, with 2013 marking the start of its validation for various clinical applications, including prostate cancer (?) (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?), traumatic brain injury (?) (?, ?, ?), heart failure (?, ?, ?, ?, ?, ?, ?, ?) and ischaemic heart disease, (?, ?, ?, ?) amongst many others. (?) To perform HP MRI, a hyperpolariser is required – a device that transiently enhances the MR signal of stable labelled isotopes by several orders of magnitude. One method to achieve this is through dissolution Dynamic Nuclear Polarisation (dDNP), which employs cryogenic temperatures, high magnetic field strengths (>3>3 T), and microwave irradiation. A typical sample for dDNP comprises a metabolic substrate mixed with a chemical radical species and optionally also a glassing agent such as glycerol. At low ($\sim 1$ K) temperature, the unpaired electron spins present in the radical become thermally polarised. When irradiated with microwaves, a population inversion of these electronic spins can be created, and they are driven to relax, primarily through dipolar interactions, and ultimately transfer their polarisation to nearby nuclei. The net process behind DNP therefore is that nuclear polarisation above that at thermal equilibrium is created. Once the desired nuclear signal enhancement is achieved, a heated and pressurised solvent is introduced to rapidly melt the solid-state sample and convert it into a form suitable for injection into a living system. Following injection, the enzymatic conversion of the hyperpolarised substrate into other metabolic intermediates can be temporally and spatially monitored in vivo through magnetic resonance experiments such as MRI.
[…]
@@
However, regardless of whether the electron source is chemical or photon-based, there are substantial technical limitations on the ability of clinical sites to ship and receive compounds for hyperpolarisation with the radicals in situ. For instance, the chemical instability of the trityl radical — the most commonly used chemical radical — renders centralised production of pre-filled fluid paths for global distribution infeasible; fluid paths are transported typically at $195$  or 253 K. (?, ?, ?) In contrast, while low-energy gamma-generated radicals are reported to be stable at room temperature for several months to produce a single sample, irradiation can take days to months, and reported nuclear enhancement factors are on the order of $10^{2}$ and not the $10^{5}$ obtained with chemical radicals. (?) DNP with UV-generated radicals is only possible if samples are optically transparent (e.g. pyruvic acid) and require low-temperature (cryogenic) transportation (?), which incurs a substantial expense and is not routinely performed as precise control of magnetic field $B$ and temperature $T$ is required over many orders of magnitude. (?) Ultimately, sites conducting human dDNP experiments generally require local access to a pharmacy equipped with sterile compounding facilities to comply with the high degree of sterility assurance required. This is complex from a regulatory view, (?) and is considered a substantial financial burden that may ultimately limit the accessibility of the technique.
@@
Here, we present an alternative approach for generating sterilised samples for clinical HP MRI that are suitable for convenient transport and centralised manufacture, potentially offering an alternative for sites that do not have access to a pharmacy. Using an ultra-high dose rate 6 MeV electron linear accelerator (linac), we demonstrate the creation of endogenous radicals within polycrystalline alanine, the concentration of which can be easily tuned by adjusting properties of the e-beam. Alanine is commonly used for radiation dosimetry, and the radicals generated upon irradiation are known to be stable for many years at room temperature if stored anhydrously. (?, ?) To improve thermal conductivity and permit efficient spin diffusion between polarising centres, irradiated polycrystalline alanine was dispersed in glycerol, resulting in radical lifetimes that were stable for several weeks when stored at room temperature with silica desiccant gel. Upon dissolution, radicals were quenched, yielding a non-toxic, (?, ?) directly injectable substrate. At 6.7 T, we were able to obtain nuclear polarisation levels of approximately 20% and discovered that the polarisation transfer is not simply explained by the solid effect or models of thermal mixing with the known stable alanine radical (SAR) $g$-factor and hyperfine couplings at this field strength. This is unlike similar analytic models of DNP with pyruvate and disperse chemical free radicals such as trityl (?). Crucially, the hyperpolarised 13C label on alanine was successfully transferred to pyruvate and lactate after injection in vivo, demonstrating the potential of this method for metabolic imaging.
@@ 2.1 Irradiation of Samples
We undertook Electron Paramagnetic Resonance (EPR) spectroscopy to quantify radical species generated by electron irradiation. As radiation damage in dry, natural-abundance (NA) alanine is known to produce stable alanine radicals species (SARs) that have been extensively studied, both via EPR and theoretically, (?, ?, ?, ?, ?, ?, ?, ?) we compared experimental EPR spectra to that of TEMPO ((2,2,6,6-Tetramethylpiperidin-1-yl)oxyl), an aminoxyl radical used for DNP, and additionally quantitatively verified their dose/response curve. Endogeneous alanine radicals had a broader linewidth than TEMPO (Fig. 1C) and, as expected an increase in the irradiation dose led to a greater spin count, saturating at $\sim$70 kGy (Fig. 1D) for the irradiated sample. We found that [1-13C]alanine had a drastically different EPR spectra compared to NA alanine, reflecting a hyperfine interaction to the 13C atom (not altering the corresponding spin count; SI §6.7). The $g$-tensor and proton/nitrogen hyperfine interaction constants $A$ for NA alanine have previously been well documented. (?, ?, ?, ?) however, no anisotropic such terms for [1-13C]alanine have been measured experimentally, although they have been predicted by Density Functional Theory calculations. (?) The EPR spectra obtained for both the NA and labelled samples at room temperature can be modelled reasonably using published hyperfine coupling values and a weighted combination of two SARs (cf. SI §6.7). For the purposes of predicting DNP behaviour, a single dominant species was considered and we modelled our system as a powder containing the labelled predominant SAR species, known as R1 (13C-SAR1). We determined using a constrained Bayesian optimisation routine (?) the values of the hyperfine coupling tensor to the 13C, $A_{13C}$, and the $g$-tensor(Fig. 1F; SI §6.8); the principal values determined for $A_{13C}$ were [1.80266,15.8427, 64.6933][-1.80266,\,-15.8427,\,64.6933] MHz with Euler angles of [1.50914, 1.30749, 0.761318][1.50914,\,1.30749,\,0.761318] rad in the previously reported molecular frame, (?) and these are a similar order of magnitude to the previously reported theoretical values calculated via Density Functional Theory (DFT), (?) although we note the accuracy of the $A_{13C}$ hyperfine values obtained is limited by the resolution of the linewidth of the EPR spectrum obtained. We additionally measured the electronic $T_{1e}$ and apparent $T_{2}$, $T_{m}$ at 5 K, and found that 100 kGy irradiated alanine had a long $T_{1}$ (7.6 s) in comparison to OX063 (1.4 s) but similar $T_{m}$ (332 ns vs. 370 ns; c.f. SI §6.7).
@@ 2.2 DNP via Irradiation-Induced Radicals
We first tested the feasibility of utilising the irradiation-induced radicals for DNP by exploring their polarisation characteristics and stability during sample storage. We investigated both dry powder mechanically secured in place with a separate snap-frozen “plug” of anhydrous glycerol and an alanine and glycerol mixture. This plug prevented the dry alanine powder from being displaced by any turbulent gas flow present, and additionally matched the dielectric loading conditions of the alanine/glycerol mixture, preserving RF coil loading and microwave power delivery between the two conditions. We further explored the comparatively limited effect of altering microwave power; c.f. SI §6.3. We found that the obtained nuclear polarisation increased with increasing radical concentration, and that substantially greater nuclear enhancement at 3.35 T was obtained with alanine crystallites dispersed in the glycerol matrix (Fig. 2A). These mixtures were stable for weeks independent of radiation dose, as quantified by EPR (Fig. 2B), with no significant change in concentration vs time (c.f. SI §6.4). The addition of glycerol did not change the electronic environment at room temperature as seen by EPR (Fig. 2C), but did at 5 K (Fig. 2D) with substantially altered hyperfine interaction constants and a shift in the mean isotropic electronic $g$-value (from $g=2.003698$ to $g=2.002862$, shown integrated in Fig. S11) that we were not able to model. The shape of the EPR spectrum at 5 K is likely indicative of a larger hyperfine interaction to the 13C atom, similar to that previously observed for the primary radical species which is only stable at low temperature; (?) indicating a further distribution over interconverting molecular species (c.f. SI §6.8). A frequency sweep at 3.35 T (Figure. S2) was difficult to interpret, with a small lobe of nuclear enhancement potentially present far away from the central maximum.
@@
Given the comparatively wide EPR linewidth of alanine radicals at 3.35 T, we elected to undertake DNP at 6.7 T. Under these conditions the relatively small degree of $g$ anisotropy would be expected to scale proportional to $B_{0}$ whereas the strong hyperfine couplings are field independent (depending primarily on the Fermi contact mechanism). We therefore expected to a first approximation an improvement in DNP efficiency through an analytic model of the thermal mixing process of polarisation growth based on the Provotorov equations, (?) as the effective width of the EPR lineshape $D$ would be expected to decrease, and the resolvable resolution of the DNP frequency sweep serve would, we believe, be predicted to increase at higher field. We additionally explored radiation doses (30 kGy, 50 kGy, and 70 kGy) to reflect the concentrations of broad line width exogenous chemical radical species typically used for DNP (i.e., 30 mM, 50 mM, and 70 mM). To our surprise, we did not observe the characteristic “bimodal” frequency sweep pattern associated with either the well-resolved solid effect or thermal mixing mechanisms, the latter of which is expected to describe DNP with nitroxide (?) or trityl (?) radicals and produces maxima approximately at the sum or difference of the nuclear and electron Larmor frequencies ($\omega_{e}\pm\omega_{N}$). Instead we clearly observed an asymmetric “hump” at higher frequency to the major lobe of the curve that decreased with increasing radiation dose (i.e. approximately at $\omega_{e}$ alone) as illustrated in Fig. 3A. At lower doses (30 kGy), we observe a bimodal microwave profile with a prominent second lobe that progressively diminishes as the radiation dose increases to 50 kGy, and ultimately disappears entirely at 70 kGy. This systematic change suggests a transition in the underlying polarisation transfer mechanism that correlates directly with radical concentration. A biexponential build-up profile was also observed (Fig. 3B) following irradiation at the maximum microwave frequency, with a fast component decreasing from 22.9 to 14.9 s and a slow component increasing from 0.92 to 6.88 hours as the radiation dose increased from 30 to 70 kGy (c.f. Table S2). This trend correlates with the disappearance of the bimodal frequency profile: samples exhibiting pronounced bimodal behavior (30 kGy) show faster overall polarisation kinetics, while those with unimodal profiles (70 kGy) exhibited slower kinetics. Whilst some degree of biexponential behaviour is expected for dilute electronic sources owing to spin diffusion, (?, ?) the longer time constant would decrease with increasing radical concentration – which we did not observe. A summary of fitted time constant parameters is provided in Table 1.
@@
To understand this difference upon the addition of glycerol, we undertook X-ray diffraction. Data were obtained from natural abundance polycrystalline alanine and polycrystalline alanine dispersed in glycerol; a spherical harmonic Rietveld crystallographic texture refinement (?) aiming to quantify any degree of order in the sample was undertaken with MAUD. (?) Using a neutron-diffraction crystal structure obtained on the SXD instrument (?) at the ISIS neutron source, (?) we demonstrated strong partial ordering of the sample (Fig. 3C) that changed (and strengthened) upon addition of glycerol (Fig. 3D). As is common in texture analysis, we express orientation in units of multiples of a random distribution (mrd), where a sample without preferred orientation is one uniformly; alanine and glycerol mixtures had orientation preferences occurring >10>10 mrd and are partially ordered. This was consistent with ab initio liquid-state molecular dynamic simulations performed on two semi-infinite glycerol/alanine domains in LAAMPs (?) with EMC (?) that predicted the immiscibility of the two substances (SI §6.10).
@@
Given the wide EPR lineshape of the stable alanine radical, we investigated the use of a frequency-swept microwave pulse sequence for DNP. This resulted in an increased polarisation transfer and increased enhancement, within a factor of 2.5 of the “gold standard” pyruvate with trityl radical (Fig. 4A). In an attempt to understand the mechanism behind this observation, we simulated the fitted EPR lineshape under DNP conditions by modelling our room temperature data as containing 13C-SAR-1. We found that a polycrystalline distribution of R1 radicals have distinct spectral density functions $g(\omega)$ depending on preferred orientation, quantified by the order parameter $\lambda$ where the angle $\alpha$ between the molecular $z$-axis and $B_{0}$ is distributed by $p(\alpha)=\exp\left(-\lambda\left(3\cos^{2}\alpha-1\right)/2\right)$, with the orientation preference found by X-ray diffraction to correspond to a combination of these parameters (Fig. 4B), whose sum could approximately trace out the enhancement curve obtained. A detailed quantum-mechanical simulation of this pulse sequence together with the known radical parameters was then undertaken in the advanced magnetic resonance simulation environment Spinach. (?) Despite including more than 50 13C spins representing an alanine supercell with detailed measurements of its proton and carbon chemical shift tensors, (?, ?) simulating the entirety of the minute-long experiment at high temporal resolution (compared to the electron Larmor frequency) and utilising the Arcus-C supercomputing cluster, (?) we were unable to reproduce this experimental result and instead predicted a bimodal curve qualitatively similar to that expected for thermal mixing (SI §6.11). Previously published analytic models of DNP via thermal mixing and the cross effect (?, ?, ?) likewise produce a predicted bimodal curve with either predicted analytic (from our model) or experimental EPR data (SI §6.12). A pseudopotential plane-wave simulation in QuantumESPRESSO (?) with our structures predicts that alanine crystals are a direct-gap semiconductor with a large (5 eV) bandgap (SI §6.13, experimentally validated (?)); were the effect of electron irradiation sufficient to dope spins into the conduction band, we would expect to see a well-resolved solid effect mechanism. This, again, did not match the data. We hypothesise that the density of unpaired electrons is insufficient to form a delocalised conduction band but rather creates localised mid-gap states. This intermediate regime, where radicals are neither fully isolated molecules (as often assumed in molecular DFT calculations used for alanine radicals (?)) nor fully delocalised carriers (as in classical solid effect DNP systems such as paramagnetic salts (?)), may explain why neither approach fully captures our experimental observations. Another possibility is the competition of multiple, competing radical species that interact and interconvert, providing a superposition of radical environments. Further work is required to assess exactly the degree remaining of molecular character in these complex, diffuse, partially ordered systems which experience significant inter-radical coupling mediated by the alanine crystal lattice and glycerol matrix, necessitating models that bridge molecular and solid-state treatments that we were unable to provide.
@@
Nevertheless, encouraged by the feasibility of polarising [1-13C]alanine via irradiation-induced radicals, our focus turned to exploiting this for biomedical imaging. Empirically optimised DNP schemes compatible with our hardware produced an estimated solid state polarisation of approximately 20% (Fig. 4C), similar to that obtained with a separate sample of [1-13C]alanine and OX063 trityl radical, both of which are comparable to the “gold standard” used clinically of pyruvate with the trityl radical AH111501, which polarised to approximately 40% (c.f. C). After dissolution, the immediate quenching in aqueous media of endogenous alanine radicals (?) resulted in an increase in the apparent liquid-state $T_{1}$ at 1.4 T (Fig. 4D), as may be expected from the removal of paramagnetic relaxation centres. (?)

### v3

# Stable electron-irradiated [1-13C]alanine radicals for clinically viable metabolic imaging with Dynamic Nuclear Polarization Short Title: Electron-irradiated radicals for DNP

@@ Abstract
Dissolution Dynamic Nuclear Polarisation (dDNP) increases the sensitivity of magnetic resonance experiments by $10^{4}-$fold, permitting isotopically-labelled molecules to be transiently visible in MRI scans. dDNP requires a source of unpaired electrons in contact with labelled nuclei, cooled to $\sim$1K, and spin-pumped into a given state by microwaves. These electrons are usually chemical radicals, requiring removal by filtration prior to injection into humans. Alternative sources, such as UV irradiation, generate lower polarisation and require cryogenic transport. We present ultra-high-dose-rate electron irradiation as a novel alternative for generating non-persistent radicals in alanine/glycerol mixtures. These are stable for months at room temperature, quench spontaneously upon dissolution, are present in dose-dependent concentrations, and generate comparable nuclear polarisation to trityl radicals used clinically (20%) through a novel mechanism. This process is inherently sterilising, permitting imaging of alanine metabolism in vivo. As well as scientific novelty, this overcomes the biggest barrier to clinically translating dDNP.
@@ 1 Introduction
Hyperpolarised Magnetic Resonance Imaging (HP MRI) is a molecular imaging technique that is predominantly used to monitor metabolic fluxes in real-time by transiently making a labelled molecular probe visible to magnetic resonance experiments such as MRI. Over the past two decades, its pre-clinical use has spanned multiple fields of basic scientific and medical research, with 2013 marking the start of its validation for various clinical applications, including prostate cancer (?), traumatic brain injury (?), and ischaemic heart disease, (?) amongst many others. (?) To perform HP MRI, a hyperpolariser is required – a device that transiently enhances the MR signal of stable labelled isotopes by several orders of magnitude. One method to achieve this is through dissolution Dynamic Nuclear Polarisation (dDNP), which employs cryogenic temperatures, high magnetic field strengths (>3>3 T), and microwave irradiation. A typical sample for dDNP comprises a metabolic substrate mixed with a chemical radical species and optionally also a glassing agent such as glycerol. At low ($\sim 1$ K) temperature, the unpaired electron spins present in the radical become thermally polarised. When irradiated with microwaves, a population inversion of these electronic spins can be created, and they are driven to relax, primarily through dipolar interactions, and ultimately transfer their polarisation to nearby nuclei. The net process behind DNP therefore is that nuclear polarisation above that at thermal equilibrium is created. Once the desired nuclear signal enhancement is achieved, a heated and pressurised solvent is introduced to rapidly melt the solid-state sample and convert it into a form suitable for injection into a living system. Following injection, the enzymatic conversion of the hyperpolarised substrate into other metabolic intermediates can be temporally and spatially monitored in vivo through magnetic resonance experiments such as MRI.
Hyperpolarised pyruvate has become the most widely used substrate in HP MRI due to its molecular utility and importance in biochemistry, and the high degree of polarisation ($50\%$) that can be obtained with dDNP. (?) However, in clinical settings, its preparation is complicated by the relative toxicity of the chemical radicals used. This requires either (1) chemical engineering to produce a soluble trityl radical that can be filtered out – the approach taken in all clinical trials to date, as the LD50 of trityl radicals is approximately 8 mmol/kg (?) and they are present in samples at concentrations above this – or (2) photon irradiation to generate transient radicals that serve as the electron source necessary for dDNP and quench rapidly upon dissolution. (?) The key advantages of the latter approach are the elimination of the filtration step, enabling faster and more efficient use of the sample, and the extended duration of the enhanced MR signal (i.e. a lengthened $T_{1}$ relaxation time), as it avoids the presence of unpaired electrons that act as relaxation centres following DNP (?). To date, both ultraviolet (?) and gamma irradiation (?) have shown potential.
@@
However, regardless of whether the electron source is chemical or photon-based, there are substantial technical limitations on the ability of clinical sites to ship and receive compounds for hyperpolarisation with the radicals in situ. For instance, the chemical instability of the trityl radical — the most commonly used chemical radical — renders centralised production of pre-filled fluid paths for global distribution infeasible; fluid paths are transported typically at $195$ K. In contrast, while low-energy gamma-generated radicals are reported to be stable at room temperature for several months to produce a single sample, irradiation can take days to months, and reported nuclear enhancement factors are on the order of $10^{2}$ and not the $10^{5}$ obtained with chemical radicals. (?) DNP with UV-generated radicals is only possible if samples are optically transparent (e.g. pyruvic acid) and require low-temperature (cryogenic) transportation (?), which incurs a significant expense and is not routinely performed as precise control of $B$ and $T$ is required over many orders of magnitude. (?) Ultimately, sites conducting human dDNP experiments generally require local access to a pharmacy equipped with sterile compounding facilities to comply with the high degree of sterility assurance required. This is complex from a regulatory view, (?) and is considered a significant financial burden that may ultimately limit the accessibility of the technique.
@@
Here, we present an alternative approach for generating sterilised samples for clinical HP MRI that are suitable for convenient transport and centralised manufacture, potentially offering an alternative for sites that do not have access to a pharmacy. Using an ultra-high dose rate 6 MeV electron linear accelerator (linac), we demonstrate the creation of endogenous radicals within polycrystalline alanine, the concentration of which can be easily tuned by adjusting properties of the e-beam. Alanine is commonly used for radiation dosimetry, and the radicals generated upon irradiation are known to be stable for many years at room temperature if stored anhydrously. To improve thermal conductivity and permit efficient spin diffusion between polarising centres, irradiated polycrystalline alanine was dispersed in glycerol, resulting in radical lifetimes that were stable for several weeks when stored at room temperature with silica desiccant gel. Upon dissolution, radicals were quenched, yielding a non-toxic, directly injectable substrate. At 6.7 T, we were able to obtain nuclear polarisation levels of approximately 20% and discovered that the polarisation transfer is not simply explained by the solid effect or models of thermal mixing with the known stable alanine radical $g$-factor and hyperfine couplings at this field strength. This is unlike similar analytic models of DNP with pyruvate and disperse chemical free radicals such as trityl (?). Crucially, the hyperpolarised 13C label on alanine was successfully transferred to pyruvate and lactate after injection in vivo, demonstrating the potential of this method for metabolic imaging.
@@ 2.1 Irradiation of Samples
We undertook EPR spectroscopy to quantify radical species generated by electron irradiation. As radiation damage in dry, natural-abundance (NA) alanine is known to produce stable alanine radicals species (SARs) that have been extensively studied, both via EPR and theoretically, (?, ?, ?, ?, ?, ?, ?, ?) we compared experimental EPR spectra to that of TEMPO ((2,2,6,6-Tetramethylpiperidin-1-yl)oxyl), an aminoxyl radical used for DNP, and additionally quantitatively verified their dose/response curve. Endogeneous alanine radicals had a broader linewidth than TEMPO (Fig. 1C) and, as expected an increase in the irradiation dose led to a greater spin count, saturating at $\sim$70 kGy (Fig. 1D). We found that [1-13C]alanine had a drastically different EPR spectra compared to NA alanine, reflecting a large hyperfine interaction (not altering the corresponding spin count; SI §6.4). The $g$-tensor and proton/nitrogen hyperfine interaction constants $A$ for NA alanine have previously been well documented; (?, ?, ?, ?) however, no such terms for [1-13C]alanine have been measured. We determined using a constrained Bayesian optimisation routine (?) the values of $A$ and $g$ for the predominant SAR species, known as R1 (Fig. 1F; SI §6.5); the principal values determined for $A$ were [1.80266,15.8427, 64.6933][-1.80266,\,-15.8427,\,64.6933] MHz with Euler angles of [1.50914, 1.30749, 0.761318][1.50914,\,1.30749,\,0.761318] rad in the previously reported molecular frame (?).
@@ 2.2 DNP via Irradiation-Induced Radicals
We first tested the feasibility of utilising the irradiation-induced radicals for DNP by exploring their polarisation characteristics and stability during sample storage. We investigated both dry powder mechanically secured in place with a separate snap-frozen “plug” of anhydrous glycerol and an alanine and glycerol mixture. We found that limiting nuclear polarisation increased with increasing radical concentration, but that significantly greater nuclear enhancement at 3.35 T was obtained with alanine crystallites dispersed in the glycerol matrix (Fig. 2A). These mixtures were stable for weeks independent of radiation dose, as quantified by EPR (Fig. 2B), with no significant change in concentration vs time. The addition of glycerol did not change the electronic environment at room temperature (Fig. 2C), but did at 5 K (Fig. 2D) with substantially altered hyperfine interaction constants and a shift in the mean isotropic electronic $g$-value (from $g=2.004100$ to $g=2.00169$, shown integrated in Fig. S7) that we were not able to model.
@@
Given the comparatively wide EPR linewidth of alanine radicals at 3.35 T, we elected to undertake DNP at 6.7 T. Under these conditions the relatively small degree of $g$ anisotropy would be expected to scale proportional to $B_{0}$ whereas the strong hyperfine couplings are field independent (depending primarily on the Fermi contact mechanism). We therefore expected to a first approximation an improvement in DNP efficiency through an analytic model of the thermal mixing process of polarisation growth based on the Provotorov equations, (?) as the effective width of the EPR lineshape $D$ would be expected to decrease. We additionally explored radiation doses (30 kGy, 50 kGy, and 70 kGy) to reflect the concentrations of broad line width exogenous chemical radical species typically used for DNP (i.e., 30 mM, 50 mM, and 70 mM). To our surprise, we did not observe the characteristic “bimodal” frequency sweep pattern associated with either the well-resolved solid effect or thermal mixing mechanisms, the latter of which is expected to describe DNP with nitroxide or trityl radicals and produces maxima approximately at the sum or difference of the nuclear and electron Larmor frequencies ($\omega_{e}\pm\omega_{N}$). Instead we observed an asymmetric “hump” at higher frequency to the major lobe of the curve that decreased with increasing radiation dose (i.e. approximately at $\omega_{e}$ alone) as illustrated in Fig. 3A. At lower doses (30 kGy), we observe a bimodal microwave profile with a prominent second lobe that progressively diminishes as the radiation dose increases to 50 kGy, and ultimately disappears entirely at 70 kGy. This systematic change suggests a transition in the underlying polarisation transfer mechanism that correlates directly with radical concentration. A biexponential build-up profile was also observed (Fig. 3B).
@@
To understand this difference upon the addition of glycerol, we undertook X-ray diffraction. Data were obtained from natural abundance polycrystalline alanine and polycrystalline alanine dispersed in glycerol; a spherical harmonic Rietveld crystallographic texture refinement (?) aiming to quantify any degree of order in the sample was undertaken with MAUD. (?) Using a neutron-diffraction crystal structure obtained on the SXD instrument (?) at the ISIS neutron source, (?) we demonstrated strong partial ordering of the sample (Fig. 3C) that changed (and strengthened) upon addition of glycerol (Fig. 3D). As is common in texture analysis, we express orientation in units of multiples of a random distribution (mrd), where a sample without preferred orientation is one uniformly; alanine and glycerol mixtures had orientation preferences occurring $10~\emph{mrd}$ and are partially ordered. This was consistent with ab initio liquid-state molecular dynamic simulations performed on two semi-infinite glycerol/alanine domains in LAAMPs (?) with EMC (?) that predicted the immiscibility of the two substances (SI §6.7).
@@
Given the wide EPR lineshape of the stable alanine radical, we investigated the use of a frequency-swept microwave pulse sequence for DNP. This resulted in an increased polarisation transfer and increased enhancement, within a factor of 2.5 of the “gold standard” pyruvate with trityl radical (Fig. 4A). In an attempt to understand the mechanism behind this observation, we simulated the fitted EPR lineshape under DNP conditions. We found that a polycrystalline distribution of R1 radicals have distinct spectral density functions $g(\omega)$ depending on preferred orientation, quantified by the order parameter $\lambda$ where the angle $\alpha$ between the molecular $z$-axis and $B_{0}$ is distributed by $p(\alpha)=\exp\left(-\lambda\left(3\cos^{2}\alpha-1\right)/2\right)$, with the orientation preference found by X-ray diffraction to correspond to a combination of these parameters (Fig. 4B). A detailed quantum-mechanical simulation of this pulse sequence together with the known radical parameters was then undertaken in the advanced magnetic resonance simulation environment Spinach. (?) Despite including more than 50 13C spins representing an alanine supercell with detailed measurements of its proton and carbon chemical shift tensors, (?, ?) simulating the entirety of the minute-long experiment at high temporal resolution (compared to the electron Larmor frequency) and utilising the Arcus-C supercomputing cluster, (?) we were unable to reproduce this experimental result and instead predicted a bimodal curve qualitatively similar to that expected for thermal mixing (SI §6.8). Previously published analytic models of DNP via thermal mixing and the cross effect (?, ?, ?) likewise produce a predicted bimodal curve with either predicted analytic or experimental EPR data (SI §6.9). A pseudopotential plane-wave simulation in QuantumESPRESSO (?) with our structures predicts that alanine crystals are a direct-gap semiconductor with a large (5 eV) bandgap (SI §6.10, experimentally validated (?)); were the effect of electron irradiation sufficient to dope spins into the conduction band, we would expect to see a well-resolved solid effect mechanism and narrow EPR data accordingly. This, again, did not match the data. We hypothesise that the density of unpaired electrons is insufficient to form a delocalised conduction band but rather creates localised mid-gap states. This intermediate regime, where radicals are neither fully isolated molecules (as often assumed in molecular DFT calculations used for alanine radicals (?)) nor fully delocalised carriers (as in classical solid effect DNP systems such as paramagnetic salts (?)), may explain why neither approach fully captures our experimental observations. The emergence of a complex and strongly orientation-dependent EPR lineshape suggests that these localised states retain molecular character in diffuse partially ordered systems while experiencing significant inter-radical coupling mediated by the alanine crystal lattice and glycerol matrix, necessitating models that bridge molecular and solid-state treatments that we were unable to provide.
@@
Nevertheless, encouraged by the feasibility of polarising [1-13C]alanine via irradiation-induced radicals, our focus turned to exploiting this for biomedical imaging. Empirically optimised DNP schemes compatible with our hardware produced an estimated solid state polarisation of approximately 20% (Fig. 4C), similar to that obtained with a separate sample of [1-13C]alanine and OX063 trityl radical, both of which are comparable to the “gold standard” used clinically of pyruvate with the trityl radical AH111501, which polarised to approximately 40%. After dissolution, the immediate quenching in aqueous media of endogenous alanine radicals resulted in an increase in the apparent liquid-state $T_{1}$ at 1.4 T (Fig. 4D), as may be expected from the removal of paramagnetic relaxation centres. (?)